# Enveloping Ellis semigroups as compactifications of transformations groups

### v2

@@ 2.4. The maps $\mathfrak{E}_{*}$ and $\tilde{\mathfrak{E}}_{*}$
(1) A group $G$ which acts ultransitively on $X$ is a dense subgroup of $({\rm S}(X),\tau_{\partial})$.
Indeed, let a set $O$ be open in $({\rm S}(X),\tau_{\partial})$ and $g\in O$. Then there are $\sigma\in\Sigma_{X}$ such that $g\mbox{{\rm St}}_{\sigma}\subset O$ and $g\mbox{{\rm St}}_{\sigma}=\{h\in G\ |\ h(x)=g(x),\ x\in\sigma\}$ is an open neighbourhood of $g$. Since $G$ acts ultransitively on $X$, there is $h\in G$ such that $h(x)=g(x)$, $x\in\sigma$. Evidently, $h\in O$.
@@
(3) Any group $G$ which acts ultransitively on $X$ is oligomorphic and, hence, Roelcke precompact (see, [44], [42]). The Roelcke precompactness of $({\rm S}(X),\tau_{\partial})$ is proved in [16] (see also [37, Example 9.14]), the Roelcke precompactness of $({\rm S}_{<\omega}(X),\tau_{\partial})$ is proved in [4].
From (1) it follows that the Roelcke compactification of a group $G$ which acts ultransitively on $X$ is isomorphic to the Roelcke compactification of $({\rm S}(X),\tau_{\partial})$.
@@
Let $G$ acts ultransitively on a discrete space $X$. Then for the action $(G,\tau_{\partial})\curvearrowright X$ items (a) and (b) of Theorem 3.4 hold and $b_{r}(G,\tau_{\partial})$ and $b_{r}{\rm S}(X)$ are topologically isomorphic semitopological semigroups.
@@
3.1.2. The maps $\mathfrak{E}_{\alpha}$ and $\tilde{\mathfrak{E}}_{\alpha}$ for an ultransitive action on a chain
For an infinite chain (linearly oredered set) $X$ let $\mbox{{\rm Aut}}(X)$ be the group of automorphisms (order-preserving bijections) of $X$. A chain $X$ is called homogeneous if the action $\mbox{{\rm Aut}}(X)\curvearrowright X$ is transitive (for any $x,y\in X$ there exists $f\in\mbox{{\rm Aut}}(X)$ such that $f(x)=y$). A chain $X$ is called 2-homogeneous if for any pairs of points $x<y$ and $x^{\prime}<y^{\prime}$ there exists $g\in\mbox{{\rm Aut}}(X)$ such that $g(x)=x^{\prime}$, $g(y)=y^{\prime}$.
@@
Secondly, the smallest linearly ordered compactification $b_{m}X$ of $X$ is zero-dimensional. In any open cover of a zero-dimensional compactum $b_{m}X$ it is possible to refine a finite cover of pairwise distinct clopen intervals no one is a subset of the other. By introducing the order on the intervals ($[a_{1},b_{1}]<[a_{2},b_{2}]$ if $a_{1}<a_{2}$), we obtain a disjoint cover of the clopen intervals $J_{1}=(\leftarrow,y_{1}),\ldots,J_{k}=(y_{k},\to)$, subtracting sequentially from the $i$ interval the union of the previous ones. Let us correct the latter cover by removing endpoints of the form $(x,0)$ from the intervals (if any), and adding them as single-point clopen intervals to the corrected interval system to obtain an open cover $\omega$ of the space $b_{m}X$. Then $\omega\wedge X$ (the trace of $\omega$ on $X$) obviously coincides with the cover from $\mathcal{U}_{X}$. Since any cover of $\mathcal{U}_{X}$ is extended to the cover of $b_{m}X$ ($u_{\theta}$ is a trace of the cover $\omega_{\theta}=[\inf,(x_{1},-1)]\cup\{(x_{1},0)\}\cup[(x_{1},1),(x_{2},-1)]\cup\{(x_{2},0)\}\cup\ldots\cup[(x_{n-1},1),(x_{n},-1)]\cup\{(x_{n},0)\}\cup[(x_{n},1),\sup]$), then $\tilde{X}=b_{m}X$ and the action $\alpha:G\times X\to X$ is extended to the action $\tilde{\alpha}:G\times b_{m}X\to b_{m}X$ (for $g\in G$
@@
Since $\mbox{{\rm St}}_{(x,j)}$, $j=-1,0,1$, for the action $\tilde{\alpha}$ coinsides with $\mbox{{\rm St}}_{x}$ for the action $\alpha$ and $\mbox{{\rm St}}_{\inf}=\mbox{{\rm St}}_{\sup}=G$, one can take arbitrary $\sigma\in\Sigma_{b_{m}X}$, where $\sigma=\sigma_{X}\cup\sigma_{\Gamma}$, $\sigma_{X}\in\Sigma_{X^{0}}(=\Sigma_{X})$, $\sigma_{\Gamma}\in\Sigma_{\Gamma}\setminus\{\inf,\sup\}$.
@@
Let $f:b_{m}X\to b_{m}X$ satisfies conditions (i)–(v). Take an arbitrary nbd $O_{f}=\bigcap\limits_{k=1}^{n}[x_{k},Of(x_{k})]$ of $f$ where, without lose of generality, $\inf<x_{1}<\ldots<x_{n}<\sup$, for any $x\in X$ only one point of the form $(x,j)$, $j=-1,0,1$, is in the set $\{x_{1},\ldots,x_{n}\}$ (use (iv)), $Of(x_{k})$, $k=1,\ldots,n$ are clopen intervals, $Of(x_{k})=Of(x_{k+1})$ if $f(x_{k})=f(x_{k+1})$ and $Of(x_{k})\cap Of(x_{k+1})=\emptyset$ if $f(x_{k})<f(x_{k+1})$, $k=1,\ldots,n-1$ (use condition (i)). Moreover, if $f(x_{k})=(y_{k},0)$ (and hence, $x_{k}\in X^{0}$), then $Of(x_{k})=\{(y_{k},0)\}$, if $f(x_{k})=(y_{k},1)$, then $Of(x_{k})=[(y_{k},1),(z_{k},-1)]$, if $f(x_{k})=(y_{k},-1)$, then $Of(x_{k})=[(z_{k},1),(y_{k},-1)]$.
@@
If $f(x,0)=(y,0),g(x,0)=(z,0)\in[\inf,(x_{1},-1)]$ or $[(x_{k},1),(x_{k+1},-1)]$, $k=1,\ldots,n-1$, or $[(x_{n},0),\sup]$, then $f(x,1)=(y,1)$, $g(x,1)=(z,1)$. The pair of points $(y,0),(y,1)$ (and $(z,0),(z,1)$) is in one element of the cover $\omega$, the pair of points $(y,0),(z,0)$ is in one element of the cover $\omega$. Hence, the pair of points $f(x,1)=(y,1)$, $g(x,1)=(z,1)$ is in one element of the cover $\omega$.
@@
Indeed, take the map $f\in bG$, $f(\inf)=\inf$, $f(x)=\sup$ otherwise. Any nbd of the map $g\in bG$, $g(\sup)=\sup$, $g(x)=\inf$ otherwise, is of the form $O_{x}^{y}=\{h\in bG\ |\ h((x,0))<(y,0)\}$, $x,y\in X$. Since $f\circ g=g$ in any nbd of $g$ of the form of $O_{x}^{y}$ there is $h\in bG$ such that $h(\inf)=\inf$, $h(x)=(y,-1)$, $\inf<x<\sup$, $h(\sup)=\sup$. Hence, $f\circ h(\inf)=\inf$, $f\circ h(x)=\sup$ otherwise and cann’t belong to the arbitary nbd of $g$ the form $O_{x}^{y}$. Thus multiplication on the left in $bG$ is not continuous.
@@
Indeed, for any $g\in\mbox{{\rm Aut}}(X)$ and a nbd $\{h\in\mbox{{\rm Aut}}(X)\ |\ h(x_{i})=g(x_{i}),\ i=1,\ldots,n\}$ of $g$, ultransitivity of the action $G\curvearrowright X$ yields that there exists $f\in G\cap\{h\in\mbox{{\rm Aut}}(X)\ |\ h(x_{i})=g(x_{i}),\ i=1,\ldots,n\}$.
@@
The action of the Thompson group $(F,\tau_{\partial})$ is ultratransitive on the chain (not continuous) $\mathbb{Q}_{r}$ of dyadic rational numbers of $(0,1)$ and every automorphism from $F$ is a piecewise-linear map that have a finite number of breakpoints [10].
@@
Another resoning (which uses Roelcke precompactness of $U({\bf H})$ [45]) to prove that the maximal equiuniformity $\mathcal{U}_{\rm S}$ on $S$ is totally bounded can be found in [26, Example 3.31].
@@
where $\mbox{{\rm pr}}_{L}$ is the orthogonal projection of ${\bf H}$ onto $L$. Without lose of generality, one can assume that $\sigma=\{x_{1},\ldots,x_{n}\}$, $0\not\in\sigma$ and ${\rm span}\ (x_{1},\ldots,x_{n})=L$.
@@
    $b_{r}U({\bf H})$ is the space of all linear operators on ${\bf H}$ of norm $\leq 1$ in the weak operator topology [45],
@@
(b) From (a) it follows that $b_{r}U({\bf H})$ is the closure of $\tilde{\imath}(U({\bf H}))$ in ${\rm B}^{\rm B}$. Since $||g(x)||=||x||$, $x\in{\rm B}$, for all $g\in U({\bf H})$, $||f(x)||\leq||x||$, $x\in{\rm B}$, for $f\in\mbox{{\rm cl}}~{}\tilde{\imath}(U({\bf H}))$ (if $||f(x)||>||x||$ for some $x\in{\rm B}$, then for the nbd $Of(x)$ of $f(x)$ such that $Of(x)\cap\{y\in{\bf H}\ |\ ||y||\leq||x||\}=\emptyset$, $[x,Of(x)]\cap\tilde{\imath}(U({\bf H}))=\emptyset$).
@@
($)C$ $\tilde{f}(cx)=2|c|||x||f\big{(}\frac{cx}{2|c|||x||}\big{)}\stackrel{{\scriptstyle\frac{c}{|c|}\leq 1,\ (2)}}{{=}}2|c|||x||\frac{c}{|c|}f\big{(}\frac{x}{2||x||}\big{)}=2c||x||f\big{(}\frac{x}{2||x||}\big{)}=c\tilde{f}(x).$
@@
Let $\tilde{f}$ be any linear operator on ${\bf H}$ of norm $\leq 1$, $f$ is its restriction to ${\rm B}$. The arbitrary nbd of $f$ is of the form $V=[x_{1},Of(x_{1})]\cap\ldots\cap[x_{n},Of(x_{n})]$, where, without loss of generality, $x_{1},\ldots,x_{n}\in{\bf H}$ are linearly independent, $Ox_{1},\ldots,Ox_{n}$ are open subsets of ${\rm B}$ in the weak topology such that ${\bf H}=L\oplus{\bf H}_{1}$ and $\mbox{{\rm pr}}^{-1}(\mbox{{\rm pr}}\ Of(x_{k}))=Of(x_{k})$, $k=1,\ldots,n$, where pr is the orthogonal projection of ${\bf H}$ to the finite dimensional subspace $L$.

### v3

@@ 2.4. The maps $\mathfrak{E}_{*}$ and $\tilde{\mathfrak{E}}_{*}$
(1) A group $G$ which acts ultratransitively on $X$ is a dense subgroup of $({\rm S}(X),\tau_{\partial})$.
Indeed, let a set $O$ be open in $({\rm S}(X),\tau_{\partial})$ and $g\in O$. Then there are $\sigma\in\Sigma_{X}$ such that $g\mbox{{\rm St}}_{\sigma}\subset O$ and $g\mbox{{\rm St}}_{\sigma}=\{h\in G\ |\ h(x)=g(x),\ x\in\sigma\}$ is an open neighbourhood of $g$. Since $G$ acts ultratransitively on $X$, there is $h\in G$ such that $h(x)=g(x)$, $x\in\sigma$. Evidently, $h\in O$.
@@
(3) Any group $G$ which acts ultratransitively on $X$ is oligomorphic and, hence, Roelcke precompact (see, [44], [42]). The Roelcke precompactness of $({\rm S}(X),\tau_{\partial})$ is proved in [16] (see also [37, Example 9.14]), the Roelcke precompactness of $({\rm S}_{<\omega}(X),\tau_{\partial})$ is proved in [4].
From (1) it follows that the Roelcke compactification of a group $G$ which acts ultratransitively on $X$ is isomorphic to the Roelcke compactification of $({\rm S}(X),\tau_{\partial})$.
@@
Let $G$ acts ultratransitively on a discrete space $X$. Then for the action $(G,\tau_{\partial})\curvearrowright X$ items (a) and (b) of Theorem 3.4 hold and $b_{r}(G,\tau_{\partial})$ and $b_{r}{\rm S}(X)$ are topologically isomorphic semitopological semigroups.
@@
3.1.2. The maps $\mathfrak{E}_{\alpha}$ and $\tilde{\mathfrak{E}}_{\alpha}$ for an ultratransitive action on a chain
For an infinite chain (linearly ordered set) $X$ let $\mbox{{\rm Aut}}(X)$ be the group of automorphisms (order-preserving bijections) of $X$. A chain $X$ is called homogeneous if the action $\mbox{{\rm Aut}}(X)\curvearrowright X$ is transitive (for any $x,y\in X$ there exists $f\in\mbox{{\rm Aut}}(X)$ such that $f(x)=y$). A chain $X$ is called 2-homogeneous if for any pairs of points $x<y$ and $x^{\prime}<y^{\prime}$ there exists $g\in\mbox{{\rm Aut}}(X)$ such that $g(x)=x^{\prime}$, $g(y)=y^{\prime}$.
@@
Secondly, the smallest linearly ordered compactification $b_{m}X$ of $X$ is zero-dimensional. In any open cover of a zero-dimensional compactum $b_{m}X$ it is possible to refine a finite cover of pairwise distinct clopen intervals no none is a subset of the other. By introducing the order on the intervals ($[a_{1},b_{1}]<[a_{2},b_{2}]$ if $a_{1}<a_{2}$), we obtain a disjoint cover of the clopen intervals $J_{1}=(\leftarrow,y_{1}),\ldots,J_{k}=(y_{k},\to)$, subtracting sequentially from the $i$ interval the union of the previous ones. Let us correct the latter cover by removing endpoints of the form $(x,0)$ from the intervals (if any), and adding them as single-point clopen intervals to the corrected interval system to obtain an open cover $\omega$ of the space $b_{m}X$. Then $\omega\wedge X$ (the trace of $\omega$ on $X$) obviously coincides with the cover from $\mathcal{U}_{X}$. Since any cover of $\mathcal{U}_{X}$ is extended to the cover of $b_{m}X$ ($u_{\theta}$ is a trace of the cover $\omega_{\theta}=[\inf,(x_{1},-1)]\cup\{(x_{1},0)\}\cup[(x_{1},1),(x_{2},-1)]\cup\{(x_{2},0)\}\cup\ldots\cup[(x_{n-1},1),(x_{n},-1)]\cup\{(x_{n},0)\}\cup[(x_{n},1),\sup]$), then $\tilde{X}=b_{m}X$ and the action $\alpha:G\times X\to X$ is extended to the action $\tilde{\alpha}:G\times b_{m}X\to b_{m}X$ (for $g\in G$
@@
Since $\mbox{{\rm St}}_{(x,j)}$, $j=-1,0,1$, for the action $\tilde{\alpha}$ coincides with $\mbox{{\rm St}}_{x}$ for the action $\alpha$ and $\mbox{{\rm St}}_{\inf}=\mbox{{\rm St}}_{\sup}=G$, one can take arbitrary $\sigma\in\Sigma_{b_{m}X}$, where $\sigma=\sigma_{X}\cup\sigma_{\Gamma}$, $\sigma_{X}\in\Sigma_{X^{0}}(=\Sigma_{X})$, $\sigma_{\Gamma}\in\Sigma_{\Gamma}\setminus\{\inf,\sup\}$.
@@
[…]
@@
[…]
@@
Indeed, take the map $f\in bG$, $f(\inf)=\inf$, $f(x)=\sup$ otherwise. Any nbd of the map $g\in bG$, $g(\sup)=\sup$, $g(x)=\inf$ otherwise, is of the form $O_{x}^{y}=\{h\in bG\ |\ h((x,0))<(y,0)\}$, $x,y\in X$. Since $f\circ g=g$ in any nbd of $g$ of the form of $O_{x}^{y}$ there is $h\in bG$ such that $h(\inf)=\inf$, $h(x)=(y,-1)$, $\inf<x<\sup$, $h(\sup)=\sup$. Hence, $f\circ h(\inf)=\inf$, $f\circ h(x)=\sup$ and can’t belong to the arbitrary nbd of $g$ the form $O_{x}^{y}$. Thus, multiplication on the left in $bG$ is not continuous.
@@
Indeed, for any $g\in\mbox{{\rm Aut}}(X)$ and a nbd $\{h\in\mbox{{\rm Aut}}(X)\ |\ h(x_{i})=g(x_{i}),\ i=1,\ldots,n\}$ of $g$, ultratransitivity of the action $G\curvearrowright X$ yields that there exists $f\in G\cap\{h\in\mbox{{\rm Aut}}(X)\ |\ h(x_{i})=g(x_{i}),\ i=1,\ldots,n\}$.
@@
The action of the Thompson group $(F,\tau_{\partial})$ is ultratransitive on the chain (not continuous) $\mathbb{Q}_{r}$ of dyadic rational numbers of (0,1)(0,1)( 0 , 1 ) and every automorphism from $F$ is a piecewise-linear map that have a finite number of breakpoints [10].
@@
Another reasoning (which uses Roelcke precompactness of $U({\bf H})$ [45]) to prove that the maximal equiuniformity $\mathcal{U}_{\rm S}$ on $S$ is totally bounded can be found in [26, Example 3.31].
@@
where $\mbox{{\rm pr}}_{L}$ is the orthogonal projection of ${\bf H}$ onto $L$. Without loss of generality, one can assume that $\sigma=\{x_{1},\ldots,x_{n}\}$, $0\not\in\sigma$ and ${\rm span}\ (x_{1},\ldots,x_{n})=L$.
@@
    $b_{r}U({\bf H})$ is the space of all linear operators on ${\bf H}$ of norm 1$\leq$ 1≤ 1 in the weak operator topology [45],
@@
[…]
@@
[…]
@@
Let $\tilde{f}$ be any linear operator on ${\bf H}$ of norm 1$\leq$ 1≤ 1, $f$ is its restriction to ${\rm B}$. The arbitrary nbd of $f$ is of the form $V=[x_{1},Of(x_{1})]\cap\ldots\cap[x_{n},Of(x_{n})]$, where, without loss of generality, $x_{1},\ldots,x_{n}\in{\bf H}$ are linearly independent, $Ox_{1},\ldots,Ox_{n}$ are open subsets of ${\rm B}$ in the weak topology such that ${\bf H}=L\oplus{\bf H}_{1}$ and $\mbox{{\rm pr}}^{-1}(\mbox{{\rm pr}}\ Of(x_{k}))=Of(x_{k})$, $k=1,\ldots,n$, where pr is the orthogonal projection of ${\bf H}$ to the finite dimensional subspace $L$.